A newly discovered Earth-like planet X is 6 lightyears away. Suppose on your 20th birthday, you were sent to Planet X on board a space probe that travels at 0.600c. How old will you be when you reach Planet X?
Answer #1

Given data:

Earth like planet is 6 lightyear away from earth.

At time of boarding, age = 20 year old

Space probe travel with speed = 0.600c

Time taken to cover the distance = \frac{distance\: }{speed }=\frac{6years*c}{0.600c}= 10 \: \: \: years

So when space probe reaches on earth like planet , the age of boy will be = 20 + 10 = 30 years old.